# Probabilistic Modeling of Antibody Kinetics Post Infection and Vaccination: A Markov Chain Approach

### v1

@@ 4.2 Conditional probabilities
The goal of considering a two-event model is that the trajectories to all states are explicit, following one branch of the graph only, and thus the conditional probability densities can be expressed in a straightforward manner because all possible paths can be easily enumerated. We present the conditional probabilities from the top “level” of Fig. 1 to the bottom.
@@
and similarly, $\text{Prob}(X_{T}=\mathcal{V})=N(\bm{r})$. These are the two states in the second “level” of the graph in Fig. 1.
Now we compute Prob$(\bm{r},T|X_{T}=\mathcal{I}^{\prime})$ and Prob$(\bm{r},T|X_{T}=\mathcal{V}^{\prime})$, the probabilities of previous (single) infection and previous (single) vaccination. These are the two states in the third “level” of the graph in Fig. 1. Since the set of previously (once) infected individuals can be partitioned by the day on which they were infected (once), we can use the law of total probability and then the definition of conditional probability to rewrite the conditional probability as
@@
The expression for Prob$(\bm{r},T|X_{T}=\mathcal{V}^{\prime})$ is analogous. We note that this differs from the corresponding conditional probability in [21] by the inclusion of the product involving $s_{\mathcal{I}^{\prime}}$; this is because in our prior work, $s_{\mathcal{I}^{\prime}}(\sigma)=1$ for all $\sigma$, as we forbid reinfections.
Next, we compute probabilities on the fourth “level” that indicate new, secondary immune events. We derive Prob$(\bm{r},T|X_{T}=\mathcal{I}\mathcal{I}^{\prime})$. Using the same approach as above, we have
@@
Finally, we want to compute probabilities like Prob$(\bm{r},T|X_{T}=\mathcal{I}^{\prime}\mathcal{I}^{\prime})$, or the fifth “level” of Fig. 1. Using the definition of conditional probability, this can be found by computing the state probabilities (done above) and probabilities of the form:
@@
where we have changed notation slightly to use $\tau=\sigma-t$, the time gap between events 1 and 2, for consistency with the personal timeline models given by Eq. (5)-(6).
@@ 5 General time-inhomogeneous model with five states
We briefly consider the pros and cons of using this general five-state framework over the two-event, 13-state framework discussed in the previous section. An advantage of the general five-state approach is that it succinctly captures antibody response and immune state simultaneously, but a disadvantage is that the corresponding conditional probability models are difficult to derive. On the other hand, the two-event, 13-state model is advantageous for the straightforward nature of its corresponding conditional probability models that are indexed by the two allowed immune events, but it is limited to short time scales or modeling non-circulating diseases because any extension to include more events becomes cumbersome and computationally impractical.
@@ 5.1 Transition probabilities
The states that we consider are naïve ($N$), newly infected ($\mathcal{I}$), newly vaccinated ($\mathcal{V}$), previously infected ($\mathcal{I}^{\prime}$), and previously vaccinated ($\mathcal{V}^{\prime}$). The graph representing the transitions between states is shown in Fig. 2. We let $s_{N}(T)$ denote the weight of the degenerate edge to $N$ from $N$, or the probability of staying naïve. The probabilities $s_{\mathcal{I}N}(T)$ and $s_{\mathcal{V}N}(T)$ weight the edges to $\mathcal{I}$ from $N$ and to $\mathcal{V}$ from $N$, indicating infection or vaccination, respectively; the other transition probabilities have analogous interpretations. The only disallowed transitions are those that go from newly infected to any state other than previously infected, and the same for new vaccination.
@@
$S(T)=\begin{bmatrix}1-s_{\mathcal{I}N}(T)-s_{\mathcal{V}N}(T)&0&0&\frac{f_{N\mathcal{I}^{\prime}}(T)}{q_{\mathcal{I}^{\prime}}(T-1)}&\frac{f_{N\mathcal{V}^{\prime}}(T)}{q_{\mathcal{V}^{\prime}}(T-1)}\\ \frac{f_{\mathcal{I}N}(T)}{q_{N}(T-1)}&0&0&\frac{f_{\mathcal{I}\mathcal{I}^{\prime}}(T)}{q_{\mathcal{I}^{\prime}}(T-1)}&\frac{f_{\mathcal{I}\mathcal{V}^{\prime}}(T)}{q_{\mathcal{V}^{\prime}}(T-1)}\\ \frac{f_{\mathcal{V}N}(T)}{q_{N}(T-1)}&0&0&\frac{f_{\mathcal{V}\mathcal{I}^{\prime}}(T)}{q_{\mathcal{I}^{\prime}}(T-1)}&\frac{f_{\mathcal{V}\mathcal{V}^{\prime}}(T)}{q_{\mathcal{V}^{\prime}}(T-1)}\\ 0&1&0&1-s_{\mathcal{I}\mathcal{I}^{\prime}}(T)-s_{\mathcal{V}\mathcal{I}^{\prime}}(T)-s_{N\mathcal{I}^{\prime}}(T)&0\\ 0&0&1&0&1-s_{\mathcal{V}\mathcal{V}^{\prime}}(T)-s_{\mathcal{I}\mathcal{V}^{\prime}}(T)-s_{N\mathcal{V}^{\prime}}(T)\end{bmatrix},$ (24)
@@
Provided the transition probabilities in Fig. 2 are nonzero, we note that every state in this Markov chain can communicate with every other, that is, it is possible to go from any state to any other state in finitely many time steps. Moreover, the period of every state is $1$, making this Markov chain irreducible and aperiodic. By the Perron–Frobenius theorem, we know that a unique stationary distribution exists, i.e. there is a unique normalized eigenvector corresponding to the eigenvalue $1$, with nonnegative entries. For the transition matrix $S(T)$ from Eq. (24), $\mathbf{e_{v}}(T)$ is the stationary distribution.
@@
We ideally want to construct a limiting distribution, i.e. a population distribution into different states in the limit of time. However, as this chain is time-inhomogeneous, whether there is a corresponding limiting distribution will additionally depend on the time-dependent incidence rates; this is an intriguing mathematical question which we are exploring in ongoing research. We will analyze the limiting distribution for a time-homogeneous Markov chain in Sections 5.3 and 7 under the assumption that the incidence rates stabilize over time.
@@ 5.2 Conditional probabilities
Here, the summand is explicitly broken up into the product of the antibody response to a particular sequence and the probability of that sequence. Let $\mathbf{\Sigma}$ be the vector of times in absolute timeline when the events happen, i.e., $\Sigma_{i}=\sum\limits_{j=1}^{i}\tau_{j}$. Then, we can explicitly calculate the probability of a given sequence of events with given $\bm{\tau}$,
@@ 5.3 Time-homogeneous model
Modifying the probability calculated in Eq. (28) by removing the dependence of time and adding the possibility of repeated return to naïve, we obtain
@@ 6 Example applied to SARS-CoV-2 antibody data
To validate our models, we use clinical data from [23] and [22] (combined because they used the same study design and measurement system), and then [24], which we will refer to as Datasets 1 and 2, respectively. The data are spike immunoglobulin G (IgG) measurements taken: before any immune event, denoted by N as naïve; after a single infection (but with no vaccination), denoted by I; after a two-dose vaccination sequence (but with no infection), denoted by V; after vaccination and a booster, denoted by VV (only for Dataset 1); and after a single vaccination dose following an infection, denoted by VI. Previously infected samples were taken from individuals whose infections were confirmed via reverse transcription polymerase chain reaction (RT-PCR) or antibody test, vaccinations were documented, and the naïve samples were collected pre-pandemic, pre-vaccination, or confirmed negative via PCR or antibody test. For Dataset 1, these values are considered together to have units of IU/mL, and for Dataset 2, the measurements are recorded in area under the curve (AUC) dimensionless units. The measurements from both datasets are log-transformed similarly to [26, 20, 27] to yield the unit-less, one-dimensional measurement
@@ 6.1 Dataset 1
Sample sizes for Dataset 1 are: N ($n=91$), I ($n=130$), V ($n=453$), VI ($n=56$), and VV ($n=275$). MLE applied to a gamma distribution for the naïve sample measurements yields optimal parameters of $\alpha_{N}=1.23$ and $\beta_{N}=0.256$. We let the exponents vary in the optimization as $k_{\mathcal{I}_{1}},k_{\mathcal{V}_{1}}$, or $k_{\mathcal{V}_{2}}$ in (5). The optimal parameter values are given in Table 1. Plotting the two-event model in two dimensions becomes difficult because there are two time variables involved, so Fig. 3 was generated using characteristic times between events of $\tau=100$ days between infection and vaccination, and Fig. 4 was generated using $\tau=275$ days for the booster visualization and $\tau=350$ days for the breakthrough infections. These are all relatively short number of days between events compared to those recorded; the corresponding videos show all times elapsed between events. Visualizations of the models without personal trajectories are shown in Figs. B1 and B2.
@@
Figure 3: Log-transformed antibody measurements from the naïve, infected (I), and infected then vaccinated (VI) populations from Dataset 1 with corresponding probability models and personal trajectories (thin, solid black lines).
Figure 4: Log-transformed antibody measurements from the vaccinated (V) and boosted (VV) populations from Dataset 1 with corresponding probability models and personal trajectories (thin, solid black lines).
For the single infection model (I) and data in Fig. 3, 66/82 (80 %) of subjects have all measurements contained within the model contours, and 49/82 (60 %) of subjects have more than half of their measurements with evaluated probabilities above 0.2, what we determine to be a reasonable “high probability” threshold (the maximum value observed is 0.84). In Fig. 4, for the single vaccination model (V) and data, 177/251 (71 %) of subjects have all measurements contained within the model probability contours (probability $>$ 0.001), and 122/251 (49 %) of subjects have more than half of their measurements with evaluated probabilities above 0.2 (the maximum value observed is 0.72).
@@ 6.2 Dataset 2
Sample sizes for Dataset 2 are: I ($n=562$), V ($n=410$), and VI ($n=261$). The log-scaled but uncorrected data from Dataset 2 are shown in Fig. 5. Notably, the characteristic decrease in antibody levels is missing for the infected samples. The AUC values from Dataset 2 are calculated using the optical density (OD) measurements from five serial dilutions (1:100, 1:300, 1:900, 1:2700, 1:8100). However, this maximal titration value may need to be raised because the OD values at a 1:8100 dilution were still nonzero and non-negligible for many samples. Therefore, the actual infection and vaccination antibody levels, especially those recorded at early times, may be higher than those reported. Vaccination response tends to lead to higher measurements than infection, and the maximal titration value was selected based on the previously infected samples, providing further evidence that vaccination levels may actually be higher than those recorded. Thus, before validating our models on Dataset 2, we performed a titration-extrapolation of all eligible measurements.
@@
To adjust reported AUC measurements in Dataset 2, we received the optical density (OD) values corresponding to five serial dilutions (1:100, 1:300, 1:900, 1:2700, 1:8100) that were used to calculate the reported AUC [24]. Dilution curves of OD plotted against increasing dilution factor that were not monotonically decreasing were excluded from the analysis and any negative OD values are not included in the AUC calculation. Some samples did not have OD measurements available. After applying our exclusion criterion, 853/1244 samples were titration-extrapolated (68.6%). The AUC is computed for the given dilutions via a trapezoidal rule. We plot the computed AUC against the reported AUC in Fig. 6a. A linear relationship is observed for most samples, which is given by
@@
This discrepancy may be due to background level subtraction step during pre-processing. Since the values seem to fall along a line of slope one with a negative intercept, the intercept of the best-fit line is recorded and used to shift the computed values to “match” the reported AUC. Then, each dilution curve is extrapolated to the next titration level (1:24300) via a quadratic fit (dashed lines in Fig. 6b), and the AUC is re-calculated using a trapezoidal rule, as long as the extrapolated OD value is nonnegative. Fig. 7 shows personal trajectories with extrapolated AUC values for all samples that met the inclusion criteria, broken down by class. On average, infection antibody levels decay more slowly than those of vaccinated individuals. We observe more of a decrease in antibody response for some infected trajectories post titration-extrapolation; many personal trajectories were fairly flat in the original data. We also observe a steeper decrease in antibody response for the vaccinated and vaccinated after infection measurements.
@@
Figure 7: Log-transformed antibody measurements from the infected (I), vaccinated (V), and infected then vaccinated (VI) populations from Dataset 2 with personal trajectories since the corresponding event. All samples’ AUC values with recorded and monotonically decreasing OD values have been titration-extrapolated using the next dilution level.
Dataset 2 did not include spike IgG measurements for naïve samples, which provide an important baseline for our models, as (4) builds off the parameters in (2). Therefore, for Dataset 2, we generate synthetic naïve data with which to build the models for previously infected and vaccinated antibody responses $t$ days after infection or vaccination, and use $\alpha_{n}=18.2$ and $\beta_{n}=0.152$. These parameters are chosen in such a way that they provide a reasonable guess for the distribution of previously infected and vaccinated responses at time zero post infection or inoculation. The optimal parameters are shown in Table 2. Plotting the two-event model becomes difficult because there are two time scales involved, so Fig. 8 was generated using $\tau=150$ days, which is a relatively short number of days between events. Visualizations of the models without personal trajectories are shown in Fig. B3.
@@
Figure 8: Log-transformed antibody measurements from the naïve (synthetic), infected (I), vaccinated (V), and infected then vaccinated (VI) populations from Dataset 2 with personal trajectories with corresponding probability models. Measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level.
In Fig. 8, for the single vaccination model (V) and data, 191/196 (97 %) of subjects have all measurements contained within the model probability contours (probability $>$ 0.001), and 131/196 (67 %) of subjects have more than half of their measurements with evaluated probabilities above 0.2 (the maximum value observed is 0.31). For the single infection model (I) and data, 302/335 (90 %) of subjects have all measurements contained within the model contours, and 70/335 (21 %) of subjects have more than half of their measurements with evaluated probabilities above 0.2 (the maximum value observed is 0.35). The vaccine response peaks between 10 and 30 days post first dose, although the data is tightly clustered in time, as is to be expected from measurements obtained via a regimented clinical trial. The infection response does not exhibit a strong peak, even less so than the vaccine response and much less in comparison to Dataset 1. This may be due to different measurement systems and data pre-processing. There are measurements with roughly similar and maximal values between 15 and 100 days post illness onset. Overall, the vaccine response is more strongly clustered in the middle of the probability model, whereas a significant number of infected measurements show a lower than average response. This is to be expected, as asymptomatic infections, varying viral load, and difficulty in defining the index date of infection create additional variability in a population’s immune response to infection. For the vaccination after infection model (VI), for those 67 subjects with a “crossover” event; i.e., with measurements belonging to both I and VI, we found that 55/67 (82 %) of these subjects have all measurements contained within the model contours (probability $>$ 0.001) and 15/67 (22 %) of these subjects have more than half of their measurements with evaluated probabilities over 0.2 (the maximum observed value is 0.36).

### v2

@@ 4.2 Conditional probabilities
The goal of considering a two-event model is that the trajectories to all states are explicit, following one branch of the graph only, and thus the conditional probability densities can be expressed in a straightforward manner because all possible paths can be easily enumerated. We present the conditional probabilities from the top “level” of Figure 1 to the bottom.
@@
and similarly, $\text{Prob}(X_{T}=\mathcal{V})=N(\bm{r})$. These are the two states in the second “level” of the graph in Figure 1.
[…]
@@
[…]
@@
[…]
@@
where we have changed notation slightly to use $\tau=\sigma-t$, the time gap between events 1 and 2, for consistency with the personal timeline models given by (5)-(6).
@@ 5 General time-inhomogeneous model with five states
We briefly consider the pros and cons of this general five-state framework versus the two-event, 13-state framework discussed in the previous section. An advantage of the general five-state approach is that it succinctly captures antibody response and immune state simultaneously; a disadvantage is that the corresponding conditional probability models are difficult to derive. On the other hand, the two-event, 13-state model is advantageous for the straightforward nature of its corresponding conditional probability models that are indexed by the two allowed immune events, but it is limited to short time scales or modeling non-circulating diseases because an extension to more events is cumbersome and computationally impractical.
@@ 5.1 Transition probabilities
The states that we consider are naïve ($N$), newly infected ($\mathcal{I}$), newly vaccinated ($\mathcal{V}$), previously infected ($\mathcal{I}^{\prime}$), and previously vaccinated ($\mathcal{V}^{\prime}$). The graph representing the transitions between states is shown in Figure 2. We let $s_{N}(T)$ denote the weight of the degenerate edge to $N$ from $N$, or the probability of staying naïve. The probabilities $s_{\mathcal{I}N}(T)$ and $s_{\mathcal{V}N}(T)$ weight the edges to $\mathcal{I}$ from $N$ and to $\mathcal{V}$ from $N$, indicating infection or vaccination, respectively; the other transition probabilities have analogous interpretations. The only disallowed transitions are those that go from newly infected to any state other than previously infected, and the same for new vaccination.
@@
$\begin{split}&S(T)\\ &=\begin{bmatrix}1-s_{\mathcal{I}N}(T)-s_{\mathcal{V}N}(T)&0&0&\frac{f_{N\mathcal{I}^{\prime}}(T)}{q_{\mathcal{I}^{\prime}}(T-1)}&\frac{f_{N\mathcal{V}^{\prime}}(T)}{q_{\mathcal{V}^{\prime}}(T-1)}\\ \frac{f_{\mathcal{I}N}(T)}{q_{N}(T-1)}&0&0&\frac{f_{\mathcal{I}\mathcal{I}^{\prime}}(T)}{q_{\mathcal{I}^{\prime}}(T-1)}&\frac{f_{\mathcal{I}\mathcal{V}^{\prime}}(T)}{q_{\mathcal{V}^{\prime}}(T-1)}\\ \frac{f_{\mathcal{V}N}(T)}{q_{N}(T-1)}&0&0&\frac{f_{\mathcal{V}\mathcal{I}^{\prime}}(T)}{q_{\mathcal{I}^{\prime}}(T-1)}&\frac{f_{\mathcal{V}\mathcal{V}^{\prime}}(T)}{q_{\mathcal{V}^{\prime}}(T-1)}\\ 0&1&0&1-s_{\mathcal{I}\mathcal{I}^{\prime}}(T)-s_{\mathcal{V}\mathcal{I}^{\prime}}(T)-s_{N\mathcal{I}^{\prime}}(T)&0\\ 0&0&1&0&1-s_{\mathcal{V}\mathcal{V}^{\prime}}(T)-s_{\mathcal{I}\mathcal{V}^{\prime}}(T)-s_{N\mathcal{V}^{\prime}}(T)\end{bmatrix},\end{split}$ (24)
@@
Provided the transition probabilities in Figure 2 are nonzero, we note that every state in this Markov chain can communicate with every other, that is, it is possible to go from any state to any other state in finitely many time steps. Moreover, the period of every state is 111, making this Markov chain irreducible and aperiodic. By the Perron–Frobenius theorem, we know that a unique stationary distribution exists, i.e. there is a unique normalized eigenvector corresponding to the eigenvalue 111, with nonnegative entries. For the transition matrix $S(T)$ from (24), $\mathbf{e_{v}}(T)$ is the stationary distribution.
@@
We ideally want to construct a limiting distribution, i.e. a population distribution into different states in the limit of time. However, as this chain is time-inhomogeneous, whether there is a corresponding limiting distribution will additionally depend on the time-dependent incidence rates; this is an intriguing mathematical question which we are exploring in ongoing research. We will analyze the limiting distribution for a time-homogeneous Markov chain in Section 5.3 and Section 7 under the assumption that the incidence rates stabilize over time.
@@ 5.2 Conditional probabilities
Here, the summand is explicitly broken up into the product of the antibody response to a particular sequence and the probability of that sequence. Let $\mathbf{\Sigma}$ be the vector of absolute timeline times when the events happen, i.e., $\Sigma_{i}=\sum\limits_{j=1}^{i}\tau_{j}$. We can explicitly calculate the probability of a given sequence of events with a given $\bm{\tau}$,
@@ 5.3 Time-homogeneous model
Modifying the probability calculated in (29) by removing the dependence of time and adding the possibility of repeated return to naïve, we obtain
@@ 6 Example applied to SARS-CoV-2 antibody data
To validate our models, we use combined clinical data from [23] and [22] and [24], which we refer to as Datasets 1 and 2, respectively. The data are spike immunoglobulin G (IgG) measurements taken: before any immune event, denoted by N as naïve; after a single infection, denoted by I; after a two-dose vaccination sequence, denoted by V; after vaccination and a booster, denoted by VV (only for Dataset 1); and after a single vaccination dose following an infection, denoted by VI. Previously infected samples were taken from individuals whose infections were confirmed via reverse transcription polymerase chain reaction (RT-PCR) or antibody test, vaccinations were documented, and the naïve samples were collected pre-pandemic, pre-vaccination, or confirmed negative via PCR or antibody test. For Dataset 1, these values are considered together to have units of IU/mL, and for Dataset 2, the measurements are recorded in area under the curve (AUC) dimensionless units. The measurements from both datasets are log-transformed similarly to [26, 20, 27] to yield the unit-less, one-dimensional measurement
@@ 6.1 Dataset 1
Measurement sample sizes for Dataset 1 are: N ($n=91$), I ($n=130$), V ($n=453$), VI ($n=56$), and VV ($n=275$). MLE applied to a gamma distribution for the naïve sample measurements yields optimal parameters of $\alpha_{N}=1.23$ and $\beta_{N}=0.256$. We let the exponents vary in the optimization as $k_{\mathcal{I}_{1}},k_{\mathcal{V}_{1}}$, or $k_{\mathcal{V}_{2}}$ in (5). The optimal parameter values are given in Table 1. Figure 3 shows the models for Dataset 1. The lightest region (yellow-ish in all plots) indicates higher probability, and the thick lines indicate probability contour levels from 0.001 (most outward) to 0.6 (most inward). The data are indicated with markers and the thin black lines indicate personal trajectories. Plotting the two-event model in two dimensions becomes difficult because there are two time variables involved, so we use characteristic times between events of $\tau=100$ days between infection and vaccination and $\tau=275$ days for the booster visualization. These are all relatively short number of days between events compared to those recorded; the corresponding Supplementary Figure (video) shows all times between vaccination events.
@@
Figure 3: Log-transformed antibody measurements from Dataset 1 with corresponding probability models and personal trajectories (thin, solid black lines). (a) naïve, infected (I), and infected then vaccinated (VI) populations; (b) vaccinated (V) and boosted (VV) populations. In (a), PDF denotes the probability distribution function.
We quantify how well the models capture the data; this will validate our modeling approach as an appropriate one. We measure this in two ways: (1) how many subjects have all measurements from their personal trajectory contained within the model contours, i.e., probability >0.001>0.001> 0.001, and (2) how many subjects have more than half of their measurements with evaluated probabilities above 0.2, what we determine to be a reasonable “high probability” threshold. The results are summarized in Table 2. For the single infection model (I) and data in Figure 3a, the maximum probability observed is 0.84, and for the single vaccination model (V) and data in Figure 3b, the maximum observed is 0.72. The results in Table 2 validate our modeling procedure, showing that a significant portion of the data lie within the contours and for the vaccination data, have high probability evaluations. This may be expected, as vaccine dosage amounts are standardized in contrast to the variable viral load of infections, which may result in a tighter distribution of vaccine responses.
Dataset Infection Vaccination
Infection then
vaccination
Vaccination,
booster
Pers. traj. in
contour lines
1 66/82 (80 %) 177/251 (71 %) 9/14 (64 %) 113/118 (96 %)
Pers. traj. eval.
meas. $\bm{>0.2}$
49/82 (60 %) 122/251 (49 %) 3/14 (21 %) 81/118 (69 %)
Pers. traj. in
contour lines
2 302/335 (90 %) 191/196 (97 %) 55/67 (82 %)
Pers. traj. eval.
meas. $\bm{>0.2}$
70/335 (21 %) 131/196 (67 %) 15/67 (22 %)
Table 2: Quantifying how well the models capture the data. Reported are how many subjects have all measurements from their personal trajectory contained within the model contours (probability >0.001>0.001> 0.001) and how many subjects have more than half of their measurements with evaluated probabilities above 0.2.
@@ 6.2 Dataset 2
Measurement sample sizes for Dataset 2 are: I ($n=562$), V ($n=410$), and VI ($n=261$). The log-scaled but uncorrected data from Dataset 2 are shown in Figure 4. Notably, the characteristic decrease in antibody levels is missing for the infected samples. The AUC values from Dataset 2 are calculated using the optical density (OD) measurements from five serial dilutions (1:100, 1:300, 1:900, 1:2700, 1:8100). However, this maximal titration value may need to be raised because the OD values at a 1:8100 dilution were still nonzero and non-negligible for many samples. Thus, the actual infection and vaccination antibody levels, especially those recorded at early times, may be higher than reported. Vaccination response tends to lead to higher measurements than infection, and the maximal titration value was selected based on the previously infected samples–further evidence that vaccination levels may actually be higher than those recorded. Thus, before validating our models on Dataset 2, we performed a titration-extrapolation of eligible measurements.
@@
To adjust reported AUC measurements in Dataset 2, we received the optical density (OD) values corresponding to five serial dilutions (1:100, 1:300, 1:900, 1:2700, 1:8100) used to calculate the reported AUC [24]. We excluded samples without OD measurements, dilution curves of OD plotted against increasing dilution factor that were not monotonically decreasing, and negative OD values. After applying our exclusion criterion, 853/1244 samples were titration-extrapolated (68.6%). The AUC is computed for the given dilutions via a trapezoidal rule. We plot the computed AUC against the reported AUC in Figure 5a. A linear relationship is observed for most samples, which is given by
@@
This discrepancy may be due to a background level subtraction step during pre-processing. Since the values seem to fall along a line of slope one with a negative intercept, the intercept of the best-fit line is recorded and used to shift the computed values to “match” the reported AUC. Then, each dilution curve is extrapolated to the next titration level (1:24300) via a quadratic fit (dashed lines in Figure 5b), and the AUC is re-calculated using a trapezoidal rule, as long as the extrapolated OD value is nonnegative. Figure 6 shows personal trajectories with extrapolated AUC values for all samples that met the inclusion criteria, broken down by class. On average, infection antibody levels decay more slowly than those of vaccinated individuals. We observe more of a decrease in antibody response for some infected trajectories post titration-extrapolation; many personal trajectories were fairly flat in the original data. We also observe a steeper decrease in antibody response for the vaccinated and vaccinated after infection measurements.
@@
Figure 6: Log-transformed antibody measurements from the infected (I), vaccinated (V), and infected then vaccinated (VI) populations from Dataset 2 with personal trajectories. All AUC values with recorded and monotonically decreasing OD values have been titration-extrapolated using the next dilution level.
Dataset 2 did not include spike IgG measurements for naïve samples, which provide an important baseline for our models, as (4) builds off the parameters in (2). Therefore, for Dataset 2, we generate synthetic naïve data with which to build the models for previously infected and vaccinated antibody responses $t$ days after infection or vaccination, and use $\alpha_{n}=18.2$ and $\beta_{n}=0.152$. These parameters are chosen in such a way that they provide a reasonable guess for the distribution of previously infected and vaccinated responses at time zero post infection or inoculation. The optimal parameters are shown in Table 3.
Figure 7 shows the models for Dataset 2. Similarly to Dataset 1, the lightest, yellow-ish region indicates higher probability, and the thick lines indicate probability contour levels from 0.001 (most outward) to 0.3 (most inward). The data are indicated with markers and the thin black lines indicate personal trajectories. This plot is visualized using a relatively short period between events, $\tau=150$ days.
The vaccine response peaks between 10 and 30 days post first dose, although the data is tightly clustered in time, as is expected from measurements obtained via a regimented clinical trial. The infection response does not exhibit a strong peak, even less so than the vaccine response and much less in comparison to Dataset 1. This may be due to different measurement systems and data pre-processing.
The vaccine response is more strongly clustered in the middle of the probability model, whereas a significant number of infected measurements show a lower than average response. This is to be expected, as asymptomatic infections, varying viral load, and difficulty in defining the index date of infection create additional variability in modeling immune response to infection.
@@
Figure 7: Log-transformed antibody measurements from the naïve (synthetic), infected (I), vaccinated (V), and infected then vaccinated (VI) populations from Dataset 2 with personal trajectories with corresponding probability models. Measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level. PDF denotes the probability distribution function.
Similarly to our analysis for Dataset 1, we quantify the goodness of fit of our modeling procedure in Table 2. For the single vaccination model (V), the maximum probability observed is 0.31, and for the single infection model (I), the maximum observed is 0.35. For the vaccination after infection model (VI), the maximum observed value is 0.36. The high percentages shown in the table validate our modeling procedure. Similar to Dataset 1, the vaccination data has especially high probability evaluations.